# FastEEC: Fast Evaluation of $N$-point Energy Correlators

### v1

@@ Abstract
Energy correlators characterize the asymptotic energy flow in scattering events produced at colliders, from which the microscopic physics of the scattering can be deduced. This view of collisions is akin to analyses of the Cosmic Microwave Background, and a range of promising phenomenological applications of energy correlators have been identified, including the study of hadronization, the deadcone effect, measuring $\alpha_{s}$ and the top quark mass. While $N$-point energy correlators are interesting to study for larger values of $N$, their evaluation is computationally intensive, scaling like $M^{N}/N!$, where $M$ is the number of particles. In this letter we develop a fast method for their evaluation, exploiting that correlations at a given angular scale are insensitive to effects at other (widely-separated) scales. For concreteness we focus on the projected energy correlator, which projects onto the largest separation between the $N$ directions. E.g. for $N=7$ we find a speed up of up to four orders of magnitude, depending on the desired accuracy. We also consider the possibility of raising the energy to a power higher than one in the energy correlator, which has been proposed to reduce soft sensitivity, and further cuts back the required computation time. These higher-power correlators are not collinear safe, but as a byproduct our approach suggests a natural method to regularize them, such that they can be described using perturbation theory. This letter is accompanied by a public code that implements our method.
@@ 1 Introduction
Energy correlators were proposed a long time ago in $e^{+}e^{-}$ collisions [1, 2, 3, 4]. They received a lot of attention in recent years due to their extension to jets [5, 6, 7]. From a physics perspective, energy correlators naturally separate effects at different scales, providing a view of e.g. the hadronization transition [8], the dead-cone effect [9] and medium effects in heavy-ion collisions [10, 11, 12, 13, 14, 15]. A range of other applications have been identified, such as the determination of the top quark mass [16, 17], where the nature of e.g. hadronization effects are different and (hopefully) better under control than for traditional observables. Indeed, energy correlators currently yield the most precise measurement at $\alpha_{s}$ from jet substructure [18]. For other recent phenomenological applications, see Refs. [19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. There have also been interesting developments on the more formal side, see Refs. [29, 30, 31, 32, 33, 34, 35, 36].
Energy correlators were first studied inside jets at hadron colliders using Open Data [8], and have been measured at STAR [37], ALICE [38] and CMS [18]. The measurement of energy correlators is made possible due to the excellent performance of detectors at the Large Hadron Collider, where the tracking system plays a crucial in accessing correlations at small angular scales. The calculation of energy correlators on tracks was developed in Refs. [6, 39, 40, 41, 42], using the track function formalism [43, 44].
In this letter we focus on the projected $N$-point energy correlator (PNEC), for which an operational definition is given by
$\frac{\mathrm{d}\sigma}{\mathrm{d}R_{L}}$ $=\int\mathrm{d}\sigma\prod_{i_{1},i_{2},\dots,i_{N}}z_{i_{1}}^{\kappa}z_{i_{2}}^{\kappa}\dots z_{i_{N}}^{\kappa}$
@@
Here $\mathrm{d}\sigma$ is the differential cross section to produce some final state, $z_{i}=p_{T,i}/p_{T,{\rm jet}}$ is the momentum fraction of particle $i$ and
$\Delta R_{ij}=\sqrt{(\Delta\eta_{ij})^{2}+(\Delta\phi_{ij})^{2}}$ (2)
is the distance between particles $i$ and $j$ in $(\eta,\phi)$ space. The delta function in Eq. (1) picks out the largest distance between the particles $i_{1},\dots i_{N}$. The default choice for the power (weight) $\kappa$ is 1, but we will also discuss other choices. In principle, one can also constrain more than just the largest distance. In that case the energy correlator would depend on up to $\tfrac{1}{2}N(N-1)$ distances, but at this point it is not clear which of these are most relevant.
Figure 1: Average time per event using Ref. [45] and our fast method, employing C/A or $k_{T}$ to cluster the jet and the indicated resolution factor $f$. While all show exponential scaling with $N$, their slopes are different leading to one or more orders of magnitude improvement at large $N$, depending on the desired accuracy.
As is clear from Eq. (1), the calculation of the $N$-point correlator for a final state with $M$ particles scales like $M^{N}/N!$, where the factorial arises because the ordering of $i_{1}$, $i_{2}$, …$i_{N}$ is irrelevant. For large values of $N$ this becomes prohibitive, and indeed studies so far have been restricted to $N\leq 6$, with $N=6$ already requiring substantial computational overhead. We address this problem here, finding that a substantial speed up is possible, depending on the desired level of accuracy. To pique your interest, we show in Fig. 1 the average time per event using our approach as function of $N$, compared to the current EnergyEnergyCorrelators package [45]. This clearly shows a speed up of multiple orders of magnitude, depending on the desired precision (which is controlled by the resolution paramter $f$ discussed later). The basic idea that we use is that correlations at a given scale are insensitive to details at much smaller scales, that don’t need to be resolved, thus reducing $M$. Furthermore, radiation separated by larger scales can be treated as independent, e.g, replacing $M^{N}$ into $m^{N}+(M-m)^{N}$, where $m$ is typically of the same order of magnitude as $M$.
@@ 2 Fast evaluation of energy correlators
The underlying idea that we utilize is as follows: for correlations at a given separation $\Delta R$ in $(\eta,\phi)$-space, radiation that is much closer together can be clustered and treated as one. The simplest way to achieve this would be to use a jet algorithm, effectively reducing the number of particles by clustering them into subjets. However, this only allows one to calculate the energy correlator at angles (much) larger than the subjet radius, $R_{L}>r$. To access correlations at small angles would require reducing the subjet radius, thereby losing the computational speed up. This can be remedied by noting that radiation separated by distances larger than $\Delta R$ can be treated independently, which our approach takes advantage of.
We use a dynamic resolution scale, such that for a given scale $\Delta R$, a subjet radius $r=\Delta R/\sqrt{f}$ is employed with $f>1$ the resolution parameter. Rather than sampling over all possible values of $\Delta R$, it is convenient to use the $\Delta R$ separations present in the particles in the jet. Concretely we achieve this by reclustering the jet using Cambridge/Aachen (C/A) [47], which results in an angular-ordered clustering tree. We then recursively traverse the tree, using the angle between two branches as $\Delta R$ and resolving the two branches using a subjet radius $r=\Delta R/\sqrt{f}$, see Fig. 2. We can then calculate the contributions to the energy correlator involving both branches, i.e. some blue and some red particles. Here the separation $\Delta R$ dominates the distance, justifying the the use of subjets as a good approximation. To also obtain the contribution involving particles from only one branch (only blue or only red) we repeat this approach recursively on each of the branches. In our implementation we use the FastJet package [48, 49] for jet reclustering and the subsequent resolution of its substructure in terms of subjets.
@@
    Recluster the final state particles into a jet using the C/A (or $k_{T}$[47, 50]. We choose a sufficiently large jet radius $R=1.5$ for this clustering step so that all the particles are inside one jet.
@@
Reclustering with the $k_{T}$ algorithm performs worse when using a constant resolution $f$: to get a similar level of accuracy requires a much larger $f$ and thus much more computing time. Interestingly, in this case using $f=k_{T,{\rm min}}^{2}$ yields rather good results within a very reasonable amount of time. We have also included this option.
Note that anti-$k_{T}$ does not improve the implementation, because anti-$k_{T}$ first clusters the energetic radiation in the jet, adding all soft radiation close to the jet boundary at the end. Consequently, declustering the branches of the first split yields many subjets, leading to very large computation times.
Additionally, our method can also be applied to compute the projected correlator where transverse momenta in Eq. (1) are weighted with power $\kappa>1$. Although, the higher power of transverse momentum weighting is collinear unsafe, it has been e.g. proposed to help mitigate the overwhelming underlying event activity in the complex environment of heavy ion collisions [15], or improve the resolution with which the top quark mass can be extracted [16]. In this case, the transverse momentum of the subjet used to calculate the energy correlator should be taken as $(\sum_{i\in\text{subjet}}p_{T,i}^{\kappa})^{1/\kappa}$ and not $\sum_{i\in\text{subjet}}p_{T,i}$, in order to agree with the full calculation of the energy correlator. Since the higher power reduces the soft sensitivity, we find that our method provides a reasonable description even with smaller values of $f$, thereby improving the computational time significantly.
@@ 3 Implementation
We start by discussing the code with C/A for $\kappa=1$. When the code is executed from the command line, the user is required to pass: the inputfile from which events should be read (input_file), the number of events (events), $N$ that specifies which point correlator to compute, the resolution factor $f$, the minimum bin value (minbin) of the histogram, the total number of bins (nbins) of the histogram and the output filename (output_file). In short, the command line syntax is:
@@
The code in which the transverse momenta of the particles are taken to a power $\kappa\neq 1$, is called eec_fast_weight and eec_fast_kt_weight. It take $\kappa>0$ as an additional command line parameter. Because this energy correlator is not collinear safe, we need to add the transverse momenta of the constituents of a subjet as $\sum_{i\in\text{subjet}}p_{T}^{\kappa}$, but we can treat the constituents as moving in the same direction, maintaining the desired speed up.
@@ 4 Results and performance
The results presented here are based on the publically available ”MIT Open Data” (MOD) [51, 52] which utilizes the reprocessed data on jets from the CMS 2011A Open Data [53, 54]. This dataset consists of jets with a transverse momenta $p_{T}\in[500,550]$ GeV and rapidity $|\eta|<1.9$. For our analysis, we have converted this dataset from MOD format to a text file format, that on each line lists: jet number, transverse momentum $p_{T}$, rapidity $\eta$ and azimuthal angle $\phi$ of a particle in that jet. This input file is also made available along with our public code and contains a total 100 000 jet events in the specified $p_{T}$ range. Note that we construct the correlator on all the final state particles in the jet and not only on charged hadrons. No detector effects or pileup subtraction is performed, as this is not the focus of our study.
@@
Figure 3: Projected $N$-point energy correlator for $N=2$ to 7, calculated using Ref. [45] and our fast method, employing C/A or $k_{T}$ to cluster the jet and the indicated resolution factor $f$. In the lower panel the relative error of these methods are shown, where we halved the number of bins to smoothen out the statistical fluctuations a bit. While the error from using $f=8$ increases from 2% to 5% as $N$ increases from 2 to 7, which can be counteracted by increasing $f$, the $k_{T}$ reclustering with $f=k_{T,{\rm min}}^{2}$ performs well across the board.
@@
Figure 4: Projected $N$-point energy correlator with $\kappa=2$ for $N=3$ (left) and $N=6$ (right), calculated using Ref. [45] and our fast method, employing C/A or $k_{T}$ to cluster the jet and the indicated resolution factor $f$. In the lower panel the relative error of these methods are shown, where we have reduced the number of bins by a factor three to smoothen out statistical fluctuations. We see that in this case the error from using $f=8$ is about 1.5% in both the cases which improves down to less than 0.5% with $f=32$. We also find that the $k_{T}$ clustering with $f=k_{T,{\rm min}}^{2}$ gives a maximum error of about 0.5% as well, though this is much smaller for the $N=3$ case.
@@
The results for our projected $N$-point correlators were obtained with a single core on a M2 chip MacBook Air both for the full calculation, for which we use the publicly available EnergyEnergyCorrelators package [45], and using our fast computation method. Fig. 1 shows the average computation time per event (in seconds) that is used for the full calculation, and each of the different choices of reclustering algorithm and resolution factor $f$ described earlier, up to the $N=7$ point correlator. We find that our method provides a substantial gain in the computation time of higher point correlators, with $N=5$ already achieving a gain by a factor of about 20, even using the slowest setting C/A $f=64$ among the choices we considered. For correlators with $N>5$, this gain is even more substantial, reaching up to a factor of about 40 for N=7 for C/A with $f=64$. The fastest option we considered, C/A with $f=8$ has a speed up of a factor 5000 for $N=7$, corresponding to an evaluation time of only about $10^{-3}s$ per event. Additionally, we observe that for the case of $k_{T}$ clustering with $f=k_{T,min}^{2}$, the algorithm provides an improvement by a factor 2 for $N=7$ when compared to C/A with $f=64$ clustering. We note that even though for $N=2$ our method does not perform as efficiently as Ref. [45] (see Fig. 1), the use of physically-motivated approximations in our approach allows for the substantial gains we observe for $N>3$.
@@
The distribution for the $N$ point correlator obtained by using different clustering schemes and choices of the resolution factor is shown in Fig. 3, along with the relative error of these methods when compared to the full computation. We show results for $N=2$ to $7$ point projected correlators, and the distributions are normalized such that the area under the curve is unity. Note that for the error plot, we halved the number of bins to reduce statistical fluctuations in the plot.
@@
There is only one region in the plot where these approximations do not perform so well, corresponding to the edge of the plot. This is not surprising, as the fine details of radiation at the jet boundary will be important to accurately describe this. However, this is not the region of interest and thus more of an artifact. Away from the jet boundary there is a power-law scaling whose exponent is related to moments of the time-like splitting functions and at very small angles there is another power-law that can be interpreted as a free-hadron gas [8].
@@
Next, we also study the behavior of correlators when higher powers of transverse momentum weights are used. Specifically, we present results for the case when $\kappa=2$, which is shown in Fig. 4 for $N=3$ and $N=6$, along with the relative error. The distributions are again normalized such that the area under the curve is unity. Note that being collinear unsafe seems to lead to much larger statistical fluctuations (the jet sample is the same as in Fig. 3). Consequently, we reduce the number of bins in the relative plot by a factor of 3, to make the trends more visible. Interestingly, we find that when the correlator is weighed by higher powers, even a small $f$ value provides a reasonable performance with an error of about 1.5% for $f=8$ which can be systematically improved to under 0.5% with $f=32$. This implies that for higher powers of $\kappa$, the computational time can be reduced by up to 4 orders of magnitude, thereby providing a substantial gain in computation time with our method.

### v2

@@ Abstract
Energy correlators characterize the asymptotic energy flow in scattering events produced at colliders, from which the microscopic physics of the scattering can be deduced. This view of collisions is akin to analyses of the Cosmic Microwave Background, and a range of promising phenomenological applications of energy correlators have been identified, including the study of hadronization, the deadcone effect, measuring $\alpha_{s}$ and the top quark mass. While $N$-point energy correlators are interesting to study for larger values of $N$, their evaluation is computationally intensive, scaling like $M^{N}/N!$, where $M$ is the number of particles. In this Letter, we develop a fast, approximate method for their evaluation, exploiting that correlations at a given angular scale are insensitive to effects at other (widely-separated) scales. This implies that the energy correlator can be computed on (sub)jets, effectively reducing $M$. Furthermore, we utilize a dynamical (sub)jet radius that allows us to obtain reliable results without restricting the angular scales being probed. For concreteness, we focus on the projected energy correlator which projects onto the largest separation between the $N$ directions. E.g. for $N=7$ we find a speed up of up to four orders of magnitude, depending on the desired accuracy. We also consider the possibility of raising the energy to a power higher than one in the energy correlator, which has been proposed to reduce soft sensitivity. These higher-power correlators are not collinear safe, but as a byproduct our approach suggests a natural method to regularize them, such that they can be described using perturbation theory. This Letter is accompanied by a public code that implements our method.
@@ 1 Introduction
Energy correlators were proposed a long time ago in $e^{+}e^{-}$ collisions [1, 2, 3, 4]. They received a lot of attention in recent years due to their extension to jets [5, 6, 7]. From a physics perspective, energy correlators naturally separate effects at different scales, providing a view of e.g. the hadronization transition [8], the dead-cone effect [9] and medium effects in heavy-ion collisions [10, 11, 12, 13, 14, 15]. A range of other applications have been identified, such as the determination of the top quark mass [16, 17], where the nature of e.g. hadronization effects are different and (hopefully) better under control than for traditional observables. Indeed, energy correlators currently yield the most precise measurement at $\alpha_{s}$ from jet substructure [18]. For other recent phenomenological applications, see Refs. [19, 20, 21, 22, 23, 24, 25, 26, 27, 28].
Energy correlators were first studied on jets at hadron colliders using Open Data [8], and have been measured at STAR [29], ALICE [30] and CMS [18]. From a theoretical perspective, these energy correlators are described by a collinear factorization formula, and can also be described in terms of the expectation value of energy flow operators [31, 32, 33, 34, 35, 36, 37]. The latter has provided the opportunity to utilize techniques developed in the context of conformal field theories to uncover the energy flow patterns in jets, resulting in numerous theoretical developments in the small angle limit [5, 6, 19, 20, 21, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Furthermore, the precision measurement of energy correlators is made possible due to the excellent performance of detectors at the Large Hadron Collider, where the tracking system plays a crucial role in accessing correlations at small angular scales. The calculation of energy correlators on tracks was developed in Refs. [6, 47, 48, 49, 50], using the track function formalism [51, 52].
In this Letter, we focus on the projected $N$-point energy correlator (PNEC), for which an operational definition is given by
$\frac{\mathrm{d}\sigma_{N}}{\mathrm{d}R_{L}}$ $=\int\mathrm{d}\sigma\prod_{i_{1},i_{2},\dots,i_{N}}z_{i_{1}}^{\kappa}z_{i_{2}}^{\kappa}\dots z_{i_{N}}^{\kappa}$
@@
Here $\mathrm{d}\sigma_{N}$ is the differential cross section to produce some final state, $z_{i}=p_{T,i}/p_{T,{\rm jet}}$ is the momentum fraction of particle $i$ and
$\Delta R_{ij}=\sqrt{(\Delta y_{ij})^{2}+(\Delta\phi_{ij})^{2}}$ (2)
is the distance between particles $i$ and $j$ in the rapidity-azimuth $(y,\phi)$ space. The delta function in Eq. (1) picks out the largest distance between the particles $i_{1},\dots i_{N}$. At the leading logarithmic accuracy, it has been shown that for the cumulative distribution defined as
$\Sigma_{N}=\int_{0}^{R_{L}}{\rm d}R_{L}^{\prime}\frac{{\rm d}\sigma_{N}}{{\rm d}R_{L}^{\prime}}\,,$ (3)
this projection to the largest distance exhibits a simple (approximate) scaling behavior $\Sigma_{N}\sim R_{L}^{\gamma(\alpha_{s},N)}$. The exponent of this scaling is related to the DGLAP anomalous dimensions, specifically the $N$-th moments of the QCD splitting functions. The default choice for the power (weight) $\kappa$ is 1, but we will also discuss other choices. In principle, one can also constrain more than just the largest distance. In that case the energy correlator would depend on up to $\tfrac{1}{2}N(N-1)$ distances, but at this point it is not clear which of these are most relevant.
Figure 1: Average time per event using Ref. [53] and our fast method, employing C/A or $k_{T}$ to recluster the jet and the indicated resolution factor $f$. While all show exponential scaling with $N$, their slopes are different leading to orders of magnitude improvement at large $N$, depending on the desired accuracy.
As is clear from Eq. (1), the calculation of the $N$-point correlator for a final state with $M$ particles scales like $M^{N}/N!$ with $M\gg N$, where the factorial arises because the ordering of $i_{1}$, $i_{2}$, …$i_{N}$ is irrelevant. For large values of $N$ this becomes prohibitive, and indeed studies so far have been restricted to $N\leq 6$, with $N=6$ already requiring substantial computational overhead [8, 50]. We address this problem here, finding that a substantial speed up is possible, depending on the desired level of accuracy. To pique your interest, we show in Fig. 1 the average time per event using our approach as function of $N$, compared to the current EnergyEnergyCorrelators package [53]. This clearly shows a speed up of multiple orders of magnitude, depending on the desired precision (which is controlled by the resolution parameter $f$ discussed later). The basic idea that we use is that correlations at a given scale are insensitive to details at much smaller scales, that don’t need to be resolved, thus reducing $M$. Furthermore, radiation separated by larger scales can be treated as independent, e.g, if the $M$ particles consist of two well-separated clusters with $m$ and $M-m$ particles, this replaces $M^{N}$ by $m^{N}+(M-m)^{N}$ for correlations at distances smaller than the separation between the two clusters.
@@ 2 Fast evaluation of energy correlators
The underlying idea that we utilize is as follows: for correlations at a given separation $\Delta R$ in $(y,\phi)$-space, radiation that is much closer together can be clustered and treated as one. The simplest way to achieve this would be to use a jet algorithm, effectively reducing the number of particles by clustering them into subjets. However, this only allows one to calculate the energy correlator at angles (much) larger than the subjet radius, $R_{L}>r$. To access correlations at small angles would require reducing the subjet radius, thereby losing the computational speed up. This can be remedied by noting that radiation separated by distances larger than $\Delta R$ can be treated independently, which our approach takes advantage of.
Figure 2: A pictorial representation of two closely spaced particles with angular orientation $\phi$.
We use a dynamic resolution scale, such that for a given scale $\Delta R$, a subjet radius $r=\Delta R/\sqrt{f}$ is employed with $f>1$ the resolution parameter. Rather than sampling over all possible values of $\Delta R$, it is convenient to use the $\Delta R$ separations present in the particles in the jet. Concretely we achieve this by reclustering the jet using Cambridge/Aachen (C/A) [55], which results in an angular-ordered clustering tree. We then recursively traverse the tree, using the angle between two branches as $\Delta R$ and resolving the two branches using a subjet radius $r=\Delta R/\sqrt{f}$, see Fig. 3. We can then calculate the contributions to the energy correlator involving both branches, i.e. some blue and some red particles. Here the separation $\Delta R$ dominates the distance, justifying the the use of subjets as a good approximation. To also obtain the contribution involving particles from only one branch (only blue or only red) we repeat this approach recursively on each of the branches. It is instructive to note that this approach can be thought of as a Taylor expansion of the exact result as follows: Consider two particles ($1$ and $2$) located at distances $R$ and $R+\delta R$ from a third particle $3$ (see Fig. 2) in the ($y-\phi$) space with an angular orientation specified by $\phi$ such that $\delta R=-\frac{R}{\sqrt{f}}\cos\phi+\frac{R}{2f}\sin^{2}\phi+{\cal O}(f^{-3/2})$. These particles are then combined to form a subjet instead of being treated separately. Concretely, this implies that for correlations at a given angular scale $R>\delta R$, the result can be approximated as
$w(R+\delta R)\approx w(R)+w^{\prime}(R)\,\delta R\,+{\cal O}[(\delta R)^{2}]\,,$ (4)
where $w(R+\delta R)$ is the energy correlator corresponding to the exact computation and $w(R)$ represents the approximate result due to the particles at $R$ and $R+\delta R$, treated as one. The leading correction in $\delta R$ that goes as $-(R/\sqrt{f})\,\cos\phi$ averages out to zero, so the extent of the relative error (due to the approximation) is governed by $1/f$. Larger values of $f$ will increase the accuracy but impact the computational time and we will show that $f=32$ already yields results at about $1-2\%$ precision with an evaluation time up-to two orders faster than the exact computation, for $N=7$. In our implementation we use the FastJet package [56, 57] for jet reclustering and the subsequent resolution of its substructure in terms of subjets.
@@
    Recluster the final state particles into a jet using the C/A (or $k_{T}$) algorithm [55, 58]. We choose a sufficiently large jet radius $R=1.5$ for this clustering step so that all the particles are inside one jet.
@@
Figure 4: To determine the scaling of $f$ with the jet $p_{T}$ for our $k_{T}$-based approach, we try two different choices: $f=k_{T,{\rm min}}^{2}/(1\,{\rm GeV})^{2}$ and $f=k_{T,{\rm min}}^{2}\times(500/p_{T})^{2}$. These agree for $p_{T}=500GeV$ but differ for other values of $p_{T}$, and we show the difference with the full result for $N=3$ (upper) and $N=4$ (lower). Clearly the second choice is the better one, with stable performance for different $p_{T}$ bins.
Reclustering with the $k_{T}$ algorithm performs worse when using a constant resolution $f$: to get a similar level of accuracy requires a much larger $f$ and thus much more computing time. Interestingly, in this case, we find that using $f=f^{\prime}k_{T,{\rm min}}^{2}$ with $f^{\prime}$ some dimensionful scale (in units of GeV-2), yields rather good results, with about a per-mille level precision for all $N$ values, within a very reasonable amount of time. When using $f=f^{\prime}k_{T,{\rm min}}^{2}$, the algorithm dynamically resolves the subsequent substructure of the splittings depending on the transverse momenta of the partons, thereby resolving more for splittings with large $k_{T,{\rm min}}$. Further, in order to determine the dependence on the jet $p_{T}$ of the dimensionful $f^{\prime}$, we perform Monte Carlo simulations and generate jets in two different $p_{T}$ ranges i.e. $p_{T}\in[500-550]\,{\rm GeV}$ and $p_{T}\in[100-150]\,{\rm GeV}$ using the anti-$k_{T}$ algorithm [59] with a radius of $R=1.0$. Using two different choices: $f=k_{T,{\rm min}}^{2}/(1{\rm GeV})^{2}$ and $f=k_{T,{\rm min}}^{2}\times(500/p_{T,{\rm jet}})^{2}$ that coincide for $p_{T,{\rm jet}}=500\,{\rm GeV}$, we study the relative error for the low $p_{T}$ bin and $N=3,4$. By comparing the relative error for two different jet $p_{T}$ ranges, from Fig. 4, we find that using $f^{\prime}\sim(500/p_{T,{\rm jet}})^{2}$ has little sensitivity across the two different $p_{T}$ ranges.
Note that anti-$k_{T}$ [59] does not improve the implementation, because anti-$k_{T}$ first clusters the energetic radiation in the jet, adding all soft radiation close to the jet boundary at the end. Consequently, declustering the branches of the first split yields many subjets, leading to very large computation times.
Additionally, our method can also be applied to compute the projected correlator where transverse momenta in Eq. (1) are weighted with power $\kappa>1$. Although, the higher power of transverse momentum weighting is collinear unsafe, it has been e.g. proposed to help mitigate the overwhelming underlying event activity in the complex environment of heavy ion collisions [15], or improve the resolution with which the top quark mass can be extracted [16]. In this case, the transverse momentum of the subjet used to calculate the energy correlator should be taken as $(\sum_{i\in\text{subjet}}p_{T,i}^{\kappa})^{1/\kappa}$ and not $\sum_{i\in\text{subjet}}p_{T,i}$, in order to agree with the full calculation of the energy correlator.
@@ 3 Implementation
We start by discussing the code with C/A for $\kappa=1$. When the code is executed from the command line, the user is required to pass: the inputfile from which events should be read (input_file), the total number of jets to be analyzed [^1] Note that we assume a jet, rather than an event, as input. (events), $N$ that specifies which point correlator to compute, the resolution factor $f$, the minimum bin value (minbin) of the histogram, the total number of bins (nbins) of the histogram and the output filename (output_file). In short, the command line syntax is:
@@
The code in which the transverse momenta of the particles are taken to a power $\kappa\neq 1$, is called eec_fast_weight and eec_fast_kt_weight. It takes $\kappa>0$ as an additional command line parameter. Because this energy correlator is not collinear safe, we need to add the transverse momenta of the constituents of a subjet as $\sum_{i\in\text{subjet}}p_{T}^{\kappa}$, but we can treat the constituents as moving in the same direction, maintaining the desired speed up.
@@ 4 Results and performance
The results presented here are based on the publicly available ”MIT Open Data” (MOD) [60, 61] which utilizes the reprocessed data on jets from the CMS 2011A Open Data [62, 63]. This dataset consists of jets with a transverse momenta $p_{T}\in[500,550]$ GeV and pseudo-rapidity $|\eta|<1.9$. For our analysis, we have converted this dataset from MOD format to a text file format, that on each line lists: jet number, transverse momentum $p_{T}$, rapidity $y$ and azimuthal angle $\phi$ of a particle in that jet. This input file is also made available along with our public code and contains a total 100 000 jet events in the specified $p_{T}$ range. Note that we construct the correlator on all the final state particles in the jet and not only on charged hadrons. No detector effects or pileup subtraction is performed, as this is not the focus of our study.
@@
Figure 5: Projected $N$-point energy correlator for $N=2$ to 7, calculated using Ref. [53] and our fast method, employing C/A or $k_{T}$ to recluster the jet and the indicated resolution factor $f$. In the lower panel the relative error of these methods are shown, where we halved the number of bins to smoothen out the statistical fluctuations a bit. While the error from using $f=8$ increases from 2% to 5% as $N$ increases from 2 to 7, which can be counteracted by increasing $f$, the $k_{T}$ reclustering with $f=k_{T,{\rm min}}^{2}$ performs well across the board.
@@
Figure 6: Projected $N$-point energy correlator with $\kappa=2$ for $N=3$ (left) and $N=6$ (right), calculated using Ref. [53] and our fast method, employing C/A or $k_{T}$ to recluster the jet and the indicated resolution factor $f$. In the lower panel the relative error of these methods are shown, where we have reduced the number of bins by a factor three to smoothen out statistical fluctuations. We see that in this case the error from using $f=8$ is about 1.5% in both the cases which improves down to less than 0.5% with $f=32$. We also find that the $k_{T}$ clustering with $f=k_{T,{\rm min}}^{2}$ gives a maximum error of about 0.5% as well, though this is much smaller for the $N=3$ case.
@@
The results for our projected $N$-point correlators were obtained with a single core on a M2 chip MacBook Air both for the full calculation, for which we use the publicly available EnergyEnergyCorrelators package [53], and using our fast computation method. Fig. 1 shows the average computation time per event (in seconds) that is used for the full calculation, and each of the different choices of reclustering algorithm and resolution factor $f$ described earlier, up to the $N=7$ point correlator. We find that our method provides a substantial gain in the computation time of higher point correlators, with $N=5$ already achieving a gain by a factor of about 20, even using the slowest setting C/A $f=64$ among the choices we considered. For correlators with $N>5$, this gain is even more substantial, reaching up to a factor of about 40 for N=7 for C/A with $f=64$. The fastest option we considered, C/A with $f=8$ has a speed up of a factor 5000 for $N=7$, corresponding to an evaluation time of only about $10^{-3}s$ per jet. Additionally, we observe that for the case of $k_{T}$ clustering with $f=k_{T,min}^{2}$, the algorithm provides an improvement by a factor 2 for $N=7$ when compared to C/A with $f=64$ clustering. We note that even though for $N=2$ our method does not perform as efficiently as Ref. [53] (see Fig. 1), the use of physically-motivated approximations in our approach allows for the substantial gains we observe for $N>3$.
@@
The distribution for the $N$ point correlator obtained by using different clustering schemes and choices of the resolution factor is shown in Fig. 5, along with the relative error of these methods when compared to the full computation. We show results for $N=2$ to $7$ point projected correlators, and the distributions are normalized such that the area under the curve is unity. For all the distribution plots shown in this Letter, we create histograms between $110^{-5}<R_{L}<$ 1 using a total of 75 logarithmic bins. Note that for the error plot, we halved the number of bins to reduce statistical fluctuations in the plot.
@@
There is only one region in the plot where these approximations do not perform so well, corresponding to the edge of the plot. This is not surprising, as the fine details of radiation at the jet boundary will be important to accurately describe this, making it sensitive to the jet algorithm. However, this is not the region of interest since in most applications of energy correlators one is interested in (much) smaller scales. One concrete application is the determination of the power-law scaling in the perturbative region, which is directly related to the moments of the time-like splitting functions. Using the CMS Open Data, it was shown in Ref. [8] that the low-point ($N=2-6$) energy correlator distribution exhibits different power law scalings as a function of the angular separation between the final state particles. In particular, away from the jet boundary, there is a perturbative power law scaling governed by splitting functions and at even smaller angles there is another power-law that can be interpreted as a free-hadron gas [8].
The approximations we make could, therefore, impact the power-law behaviour (in the perturbative region) extracted from LHC data. However, as is clear from Fig. 5, our method can still be used to reliably extract this power-law scaling at the few percent-level or less with a reasonable computation time. This implies that our method provides access to this scaling for much larger values of $N$ than currently realized through the existing EnergyEnergyCorrelators package due to the computation time.
@@
Next, we also study the behavior of correlators when higher powers of transverse momentum weights are used. Specifically, we present results for the case when $\kappa=2$, which is shown in Fig. 6 for $N=3$ and $N=6$, along with the relative error. The distributions are again normalized such that the area under the curve is unity. Note that being collinear unsafe seems to lead to much larger statistical fluctuations (the jet sample is the same as in Fig. 5). Consequently, we reduce the number of bins in the relative plot by a factor of 3, to make the trends more visible.